Mr. Balter wanted to make a plywood rectangular table top that would be 3 feet longer than it was wide. He found that by making the top 1 foot narrower and 2 feet shorter than planned, he could save $2.60 on the plywood. If the plywood he was using cost 20 cents per square foot, what were Mr. Balter's original dimensions for his table top?
Answer by checkley71(8403)

SEEING AS THE WOOD COST $.20 PER SQ FOOT THEN HE SAVED A TOTAL OF 13 SQ FEET WITH THE REDUCED TABLE SIZE THEREFORE
X(X+3)=(X-1)(X+1)+13 THIS COMPARES THE ORIGINAL SIZE X WIDE AND (X+3) LONG &
THE REDUCED SIZE OF (X-1) FOR THE WIDTH AND (X+3-2) FOR THE LENGTH
THUS WE HAVE X~2+3X=X~2-1+13 OR CANCELLING OUT THE X~2 WE ARE LEFT WITH
3X=12 OR X=12/3 OR X=4
PROOF THE ORIGINAL TABLE WAS 4 BY 7 OR 28 SQ FEET @ .20 SQ FOOT IT COST $5.60
THE REDUCED TABLE OF (4-1) AND (7-2) OR 3 BY 5 OR 15 SQ FEET @$.20 SQ FOOT IT COST $3.00 THERFORE THE SAVINGS WAS $5.60-$3.00=$2.60